# Masked Image Contrastive Learning for Efficient Visual Conceptual Pre-training

### v1

@@ Abstract
This paper proposes a scalable and straightforward pre-training paradigm for efficient visual conceptual representation called masked image contrastive learning (MiCL). Our MiCL approach is simple: we randomly mask patches to generate different views within an image and contrast them among a mini-batch of images. The core idea behind MiCL consists of two designs. First, masked tokens have the potential to significantly diminish the conceptual redundancy inherent in images, and create distinct views with substantial fine-grained differences on the semantic concept level instead of the instance level. Second, contrastive learning is adept at extracting high-level semantic conceptual features during the pre-training, circumventing the high-frequency interference and additional costs associated with image reconstruction. Importantly, MiCL learns highly semantic conceptual representations efficiently without relying on hand-crafted data augmentations or additional auxiliary modules. Empirically, MiCL demonstrates high scalability with Vision Transformers, as the ViT-L/16 can complete pre-training in 133 hours using only 4 A100 GPUs, achieving 85.8% accuracy in downstream fine-tuning tasks.
@@ 1 Introduction
Self-supervised learning (SSL) is considered the cornerstone towards building a world model, particularly in the pre-training of vision model [4, 12, 2, 19, 23, 28], attributed to its ability to acquire versatile visual representations without the need for human annotations. Currently, two dominant learning paradigms in visual self-supervised learning are the Masked Image Modeling (MIM) [12, 23, 17, 31, 11] and Contrastive Learning (CL) [4], exhibiting promising scalability characteristics for vision models, notably Vision Transformers (ViTs) [8].
Despite the success of these methods, both of these prevailing paradigms suffer from efficient visual representation, due to image sparsity and conceptual redundancy. The image sparsity leads to an excessive focus on local details during pixel-level reconstruction in MIM, rather than the highly semantical concepts. On the flip side, conceptual redundancies typically result in transformed images that fail to exhibit significant distinctions in CL. Hence, a pertinent question emerges: beyond the existing MIM and CL paradigms, how to reconcile the divergence between efficient visual representation and effective conceptual pre-training?
In elucidating this query, our initial step involves a revisit to the present pre-training paradigms, namely MIM and CL. Regarding MIM, it aims to learn visual representatives by reconstructing the masked image patches, as illustrated in Fig.1(b). Among these methods, BEiT [2] and MAE [12] are representatives. It is important to mention, that MAE indicates that there remains significant semantic redundancy within images, that only a small high-level understanding of parts, objects, and scenes is required to recover missing patches from neighbouring patches. However, pixel-level restoration is overly fine-grained for the pre-training of a vision model, excessively focusing on high frequencies and local details of the image. It runs counter to the core of pre-training, which is targeted at encapsulating high-hierarchical semantic image concepts. Despite this intricate task aids in the acquisition of visual representatives by the model, it regrettably fails to address concerns regarding pre-training efficiency.
In terms of CL, the underlying principle of it hinges on a quite simple concept: maximum agreement between varying views from a single image, as exhibited in Fig.1(a), where SimCLR [4, 5], MoCo v3 [6] and DINO [3] are typical representatives. Within this cohort, sophisticated pre-processing techniques and intricate auxiliary networks are required to capture distinct views of the image, thereby presenting a challenging task of optimizing the agreement among different views. Consequently, it can be derived that the essence of contrastive learning lies in the creation of distinct views characterized by substantial disparities. However, this endeavour presents a formidable challenge owing to the inherent conceptual redundancy prevalent in images. Moreover, large batch sizes are prerequisites for the generation of negative sample pairs, leading to a long training time and huge computing resources consumption.
@@
Driven by this analysis, we found that these two paradigms can complement each other: masked tokens have the potential to significantly diminish the conceptual redundancy inherent in images, whereas contrastive learning is adept at extracting high-level semantic features during the pre-training phase. Thus, we present a novel and straightforward paradigm for self-supervised visual representation learning: masked image contrastive learning (MiCL). MiCL tackles the above issues systematically: I) Masked image tokens offer diverse views of a single image with substantial fine-grained conceptual differences. II) Contrastive learning enables pre-training to concentrate exclusively on the high-level semantic information contained within images while disregarding high-frequency redundancies. III) The proposed paradigm obviates the need for auxiliary modules and expedites the efficient extraction of model features.
(a) Contrastive Learning
(b) Masked Image Modeling
(c) Our Masked Image Contrastive Learning
@@
Figure 1: Comparison between different pre-training paradigms. The Model in blue is the pre-training model, and the orange modules indicate auxiliary modules. (a) Contrastive Learning endeavours to maximize the agreement between different views of an image. (b) Masked Image Modeling aims to restore masked image patches. (c) Our Masked Image Contrastive Learning: Through non-overlapping masking, distinct tokens within an image are categorized as intraclass, while across-images tokens within a batch are viewed as interclass. Our objective is to enhance intraclass compactness and interclass separability through a contrastive learning approach. It is worth noting that our method does not necessitate any auxiliary modules, ensuring that no training resources are expended on redundant networks.
@@
MiCL has a particularly simple and straightforward workflow, as presented in Fig.1(c). Here is how it works: Firstly, we mask a batch of images with a high rate, dividing visible patches within one image into two non-overlapping groups. In succession, the pre-train model extracts the features of these two groups of batch image tokens, respectively. Subsequently, contrastive learning is employed to predict the correct pairings for a batch of visible image tokens. Positive samples are different visible tokens in the same image, while negative samples are from different images of the mini-batch. Finally, inspired by T-distributed classifier [26, 27], we introduce the T-distributed spherical loss to constrains the inter-class margins in the pre-training. Comprehensive experiments demonstrate the scalability and efficacy of our approaches, where ViT-L/16 can complete pre-training in 133 hours using only 4 A100 GPUs and attain an 85.8% top-1 accuracy in fine-tuning classification. In particular, our model stands out from other pre-training methods as it operates without the need for auxiliary modules or hand-crafted data augmentation to generate diverse views.
@@
    We endeavour to explore an alternative of using masked images to create diverse views with fine-grained conceptual differences for contrastive learning. By forgoing the conventional approach of employing instance-level hand-crafted data augmentation to generate distinct views, MiCL diminishes the conceptual redundancy inherent in images and improves efficiency.
@@
    Our approach eschews the reconstruction of masked images in favor of leveraging contrastive loss to steer the entire model. Independently of additional auxiliary modules, MiCL is adept at extracting high-level semantic concept features from images more efficiently.
@@
    Extensive experiments are conducted to verify the efficiency and scaling capability of our method. ViT-L/16 can complete pre-training in 133 hours using only 4 A100 GPUs with 85.8% accuracy in fine-tuning. Additionally, we have structured ablation experiments to delve into the implications of different configurations within MiCL, with a particular focus on the need of the MLP head in contrastive learning.
2 Related Works
2.1 Masked Image Modeling
Inspired by masked language modeling in NLP, the core of masked image modeling is to predict the masked part of the input image. Among them, BEiT [2] tokenizes image patches through the reconstruction of the individual image using dVAE, and then predicts the tokens of the masked patches to learn visual representation.
Similarly, MAE [12] utilizes a high masked ratio (75%) to corrupt the image and directly reconstruct the pixel-level masked image patches. Subsequently, numerous studies have referenced this paradigm for pre-training endeavours, including DropPos [23], U-MAE [31] and CAE [7]. DropPos [23] incorporates position reconstruction to bolster the spatial awareness of ViTs. U-MAE [31] introduces a uniformity loss as a regularization to the MAE loss to further encourage the feature consistency of the pre-training, and addresses the dimensional feature collapse. CAE [7] decouples the learning processes for image representation and pretext tasks, enabling the pre-trained model to prioritize image representation while disregarding the pretext task.
Additionally, following the successful implementation of MAE, its applicability has been extended to diverse disciplines, such as SiamMAE [11] and MR-MAE [9].
2.2 Contrastive Learning
Within this cohort, SimCLR [4] demands sophisticated pre-processing techniques to capture distinct views of the image, intending to present a challenging task to facilitate the acquisition of effective visual representations in pre-training. Meanwhile, large batch size is the prerequisite for the generation of negative sample pairs. Thereby, it requires a long training time and huge computing resources. Besides, DINO [3] utilizes student-teacher architecture to extract visual representation from different views. Moreover, MoCo v3 [6] applies momentum update to drive the auxiliary network. Consequently, it can be derived that the essence of contrastive learning lies in the creation of distinct views characterized by substantial disparities. However, this endeavour presents a formidable challenge owing to the inherent semantic redundancy prevalent in images. Besides, ConCL [24] generates distinct concepts through image cropping and leverages contrastive learning within a teacher-student framework to pre-train pathological images.
2.3 Combination between MIM and CL
There are also a lot of efforts to bridge the gap between mask image modeling and contrastive learning, where the integration of teacher-student models emerges as the prevailing approach. iBOT [33] employs the teacher-student network to independently encode the two augmented views, with the student network processing masked images. The objectives of MIM and CL are jointly trained for self-distillation. Likewise, MST [18] introduces a masked token strategy leveraging multi-head self-attention maps, which selectively mask the tokens of the student network based on the output self-attention map of the teacher network, ensuring vital foreground remains intact. Similarly, SiameseIM [21] employs a Siamese network featuring two branches. The online branch encodes the initial view and predicts the representation of the second view based on their relative positions. Meanwhile, the target branch generates the target by encoding the second view.
ccMIM [32] leverages a contrastive loss to aid the reconstruction task as a regularizer, facilitating the extraction of image-wide global information from both masked and unmasked patches. Likewise, ConMIM [29] produces simple intra-image inter-patch contrastive constraints as the sole learning objectives for masked patch prediction, and strengthens the denoising mechanism with asymmetric designs to improve the network pre-training. Similarly, I-JEPA [1] predicts the feature encoding of the contextual region via MIM, and employs contrastive learning to align the features of neighbouring regions at the feature level. Additionally, CoMAE [25] also applies CL to assist cross-modal MIM tasks. Besides, LGP [15] integrates MIM and CL in a sequential cascade manner: early layers are first trained under one MIM loss, on top of which latter layers continue to be trained under another CL loss.
@@ 3.1 Architecture
3.2 Masked Image for Conceptual Pre-training
@@
We apply the masking strategy of random sampling, the same as the MAE, which samples random patches without replacement following the uniform distribution. Beyond that, a low masking ratio (e.g. 30%) is implemented overall, but leads to a higher masking ratio (e.g. 65%) for individual contrastive branches, as illustrated in Fig. 2. The overarching low global masking rate facilitates the model in capturing the holistic information of the image and the interconnections among various concepts within a mini-batch. On the flip side, in each forward propagation, the model is exposed to a limited portion of the visible tokens, thereby generating diverse views with different fine-grained semantic concepts. The high masking ratio for each contrastive branch creates a pretext task that cannot be easily solved by extrapolation from simple image deformation, thus diminishing conceptual redundancy within the image, and emphasizing the localized features of the image. Furthermore, the masking strategy enables training solely on a few portions of the image, enhancing the scalability and efficiency of model training while accommodating the demands associated with large batch sizes in contrastive learning.
In terms of partitioning visible patches into two non-overlapping distinct groups, we simply employ a random sampling strategy akin to the aforementioned masking approach to mitigate potential biases, ensuring that the central positions of visible patches within the two groups remain consistent. Meanwhile, non-overlapping patches present a challenging scenario, impeding the model from relying solely on analogous patches for inference.
@@ 3.3 Efficient Contrastive Learning
We randomly select a mini-batch of $B$ instances and establish the contrastive prediction task on the visible token pairs extracted from this mini-batch, yielding a total of $2\times B$ data points. We designate feature tokens stemming from disparate groups within the same image as positive pairs. Conversely, tokens from distinct images within the same mini-batch are regarded as negative pairs. We concurrently maximise the similarity of $B$ positive pairs while minimizing the similarity of $B^{2}-B$ negative examples to drive the network.
@@
$\mathcal{L}(y_{i}^{U},y_{j}^{L})$ (3)
$=$ $-\log{\frac{\exp{(\text{sim}_{tsp}(y_{i}^{U},y_{j}^{L})\times\tau)}}{\sum_{k=1}^{2B}{\mathbbm{1}_{\left[k\neq i\right]}\exp{(\text{sim}_{tsp}(y_{i}^{U},y_{k}^{L})\times\tau)}}}}$
where $\mathbbm{1}_{\left[k\neq j\right]}\in\{0,1\}$ is an indicator function evaluating to 1 if $k\neq i$. Finally, inspired by CLIP [19], we optimize a symmetric loss over these similarity scores within a mini-batch exhibited in Algorithm 1.
@@ 3.4 Simple implementation
Algorithm 1 Pytorch-like pseudo code for the core of an implement of MiCL
The implementation of our MiCL pre-training is efficient and involves minimal specialized operations. As pseudo code depicted in Algorithm 1, we make only minor modifications based on the MAE code, mainly involving the process subsequent to the acquisition of image feature embeddings from the encoder. First, we randomly mask a subset of embedded patch tokens with a low masking ratio. In succession, listed tokens are shuffled randomly and divided into two non-overlapping groups. Following MAE [12], within positional and [CLS] embeddings, lists of tokens are encoded by the ViT. It is noteworthy that we obtain the encoded [CLS] token from the ViT directly, without the incorporation of additional auxiliary modules, even a linear head or lightweight decoder.
@@ 4 Experiments
ViT-B/16 [8] 12 768 12 86M
ViT-L/16 [8] 24 1024 16 304M
ViT-h/16 [8] 32 1280 16 632M
Table 1: Configurations of Vision Transformer models in our experiments. Block denotes the number of transformer blocks, with dim representing the dimension of the input/output channel dimension of all blocks. Heads are the number of heads in multi-head attention modules. And we also provide the parameter sizes of different models.
We conduct self-supervised pre-training on the ImageNet-1K [20] dataset with the resolution of 224$\times$224. By default, ViT-B/16 and ViT-L/16 [8] are leveraged as the backbone architecture with 800 epochs for pre-training and 40 epochs for warm-up. ViT-B/16 applies the overall masked ratio of 0.3, while ViT-L/16 is set to 0.4. And the initial base learning rate is $1.5\times 10^{-4}$. Similar to other contrastive learning [4, 19], our method relies on the large effective batch size, specifically set at 9,600 for ViT-B/16 and 2,048 for ViT-L/16. Besides, $\kappa$ is set to 64 in the T-SP metric to compute the similarity by default. Our implementation is based on MAE [12]. Details can be found in Supplementary Material.
In terms of supervised validation, MiCL is evaluated through end-to-end fine-tuning and linear probing on the ImageNet-1k dataset for classification with 100 epochs for ViT-B/16 and 50 for ViT-L/16, following common practices [3, 12, 23]. Top-1 accuracy is utilized to verify the performance of different methods.

### v2

# One Leaf Reveals the Season: Occlusion-Based Contrastive Learning with Semantic-Aware Views for Efficient Visual Representation

@@ Abstract
This paper proposes a scalable and straightforward pre-training paradigm for efficient visual conceptual representation called occluded image contrastive learning (OCL). Our OCL approach is simple: we randomly mask patches to generate different views within an image and contrast them among a mini-batch of images. The core idea behind OCL consists of two designs. First, masked tokens have the potential to significantly diminish the conceptual redundancy inherent in images, and create distinct views with substantial fine-grained differences on the semantic concept level instead of the instance level. Second, contrastive learning is adept at extracting high-level semantic conceptual features during the pre-training, circumventing the high-frequency interference and additional costs associated with image reconstruction. Importantly, OCL learns highly semantic conceptual representations efficiently without relying on hand-crafted data augmentations or additional auxiliary modules. Empirically, OCL demonstrates high scalability with Vision Transformers, as the ViT-L/16 can complete pre-training in 133 hours using only 4 A100 GPUs, achieving 85.8% accuracy in downstream fine-tuning tasks. Code is available at https://anonymous.4open.science/r/OLRS/.
@@ 1 Introduction
Self-supervised learning (SSL) is a key approach for building world models, especially for pre-training vision models [1, 2, 3, 4, 5, 6]. Its strength lies in learning versatile visual representations without relying on human annotations. Currently, the two main paradigms in visual SSL are Masked Image Modeling (MIM) [2, 5, 7, 8, 9] and Contrastive Learning (CL) [1]. Both have shown strong scalability, particularly for Vision Transformers (ViTs) [10].
Despite the success of existing methods, both MIM and CL struggle to achieve efficient visual representation. MIM focuses heavily on pixel-level reconstruction, which often prioritizes local details over high-level semantic concepts. Similarly, CL suffers from conceptual redundancy, where transformed images may lack meaningful differences. This raises a critical question: How can we bridge the gap between efficient visual representation and effective conceptual pre-training, overcoming the limitations of current MIM and CL approaches?
To address this question, we first revisit the two main pre-training paradigms: MIM and CL. MIM learns visual representations by reconstructing masked image patches (see Fig.1(b)). Notable examples include BEiT [3] and MAE [2]. MAE highlights that images contain significant semantic redundancy, meaning only a basic understanding of objects and scenes is needed to predict missing patches from their surroundings. However, pixel-level reconstruction is too detailed for pre-training vision models. It focuses excessively on high-frequency details and local features, conflicting with the goal of pre-training: learning high-level semantic concepts. While this detailed task helps models learn visual representations, it comes at the cost of pre-training efficiency.
In CL, the core idea is simple: maximize agreement between different views of the same image (see Fig.1(a)). Popular methods like SimCLR [1, 11], MoCo v3 [12], and DINO [13] use complex pre-processing and auxiliary networks to create distinct views of an image. The main challenge lies in optimizing the agreement between these views. For CL to work well, distinct views are essential. However, this is difficult because images often have conceptual redundancy. Additionally, the dependency on large-scale batches to generate sufficiently diverse negative samples imposes stringent computational constraints, notably escalating computational and time consumption overhead during training.
@@
Driven by this analysis, we found that these two paradigms can complement each other: masked tokens have the potential to significantly diminish the conceptual redundancy inherent in images, whereas contrastive learning is adept at extracting high-level semantic features during the pre-training phase. Thus, we present a novel and straightforward paradigm for self-supervised visual representation learning: occluded image contrastive learning (OCL). OCL tackles the above issues systematically: I) Masked image tokens offer diverse views of a single image with substantial fine-grained conceptual differences. II) Contrastive learning enables pre-training to concentrate exclusively on the high-level semantic information contained within images while disregarding high-frequency redundancies. III) The proposed paradigm obviates the need for auxiliary modules and expedites the efficient extraction of model features.
(a) CL
(b) MIM
(c) Our occluded image contrastive learning
@@
Figure 1: Comparison between different pre-training paradigms. The Model in blue is the pre-training model, and the orange modules indicate auxiliary modules. (a) Contrastive Learning (CL) endeavours to maximize the agreement between different views of an image. (b) Masked Image Modeling (MIM) aims to restore masked image patches. (c) Our occluded image contrastive learning: Through non-overlapping occluding, distinct tokens within an image are categorized as intraclass, while across-image tokens within a batch are viewed as interclass. Our objective is to enhance intraclass compactness and interclass separability through a contrastive learning approach. Just as a single leaf can tell the coming of autumn, we believe that a small area of the image contains the majority of the meaning of the entire image.
@@
OCL has a particularly simple and straightforward workflow, as presented in Fig.1(c). Here is how it works: Firstly, we mask a batch of images with a high rate, dividing visible patches within one image into two non-overlapping groups. In succession, the pre-train model extracts the features of these two groups of batch image tokens, respectively. Subsequently, contrastive learning is employed to predict the correct pairings for a batch of visible image tokens. Positive samples are different visible tokens in the same image, while negative samples are from different images of the mini-batch. Finally, inspired by T-distributed classifier [14, 15], we use the T-distributed spherical loss to constrain the inter-class margins in the pre-training. Comprehensive experiments demonstrate the scalability and efficacy of our approaches, where ViT-L/16 can complete pre-training in 133 hours using only 4 A100 GPUs and attain an 85.8% top-1 accuracy in fine-tuning classification. In particular, our model stands out from other pre-training methods as it operates without the need for auxiliary modules or hand-crafted data augmentation to generate diverse views.
@@
    We endeavour to explore an alternative of using masked images to create diverse views with fine-grained conceptual differences for contrastive learning. By forgoing the conventional approach of employing instance-level hand-crafted data augmentation to generate distinct views, OCL diminishes the conceptual redundancy inherent in images and improves efficiency.
@@
    Our approach eschews the reconstruction of masked images in favour of leveraging contrastive loss to steer the entire model. Independently of additional auxiliary modules, OCL is adept at extracting high-level semantic concept features from images more efficiently.
@@
    Extensive experiments are conducted to verify the efficiency and scaling capability of our method. ViT-L/16 can complete pre-training in 133 hours using only 4 A100 GPUs with 85.8% accuracy in fine-tuning. Additionally, we have structured ablation experiments to delve into the implications of different configurations within OCL, with a particular focus on the need of the MLP head in contrastive learning.
@@ 2.1 Architecture
2.2 Occluded Image for Conceptual Pre-training
@@
We apply the masking strategy of random sampling, the same as the MAE, which samples random patches without replacement following the uniform distribution. Beyond that, a low masking ratio (e.g. 30%) is implemented overall, but leads to a higher masking ratio (e.g. 65%) for individual contrastive branches, as illustrated in Fig. 2. The low global masking rate helps the model capture the overall image structure and the relationships between different concepts within a mini-batch. During each forward pass, the model only sees a small portion of the visible tokens. This creates diverse views with fine-grained semantic differences, making the task more challenging and reducing the risk of the model relying on simple patterns or redundant information. The high masking ratio for each contrastive branch ensures that the task cannot be solved by simply extrapolating from basic image transformations. This reduces conceptual redundancy and forces the model to focus on localized features. Additionally, by training on only a few parts of the image, the masking strategy also improves scalability and efficiency, making it easier to handle the large batch sizes required for contrastive learning.
To divide visible patches into two separate, non-overlapping groups, we use a random sampling strategy similar to the masking approach described earlier. This helps avoid potential biases and ensures that the central positions of the visible patches are consistent across both groups. Meanwhile, non-overlapping patches present a challenging scenario, impeding the model from relying solely on analogous patches for inference.
@@ 2.3 Efficient Contrastive Learning
We randomly select a mini-batch of $B$ instances and establish the contrastive prediction task on the visible token pairs extracted from this mini-batch, yielding a total of $2\times B$ data points. Feature tokens from different groups within the same image are treated as positive pairs, while tokens from different images in the same mini-batch are treated as negative pairs. We concurrently maximise the similarity of $B$ positive pairs while minimizing the similarity of $B^{2}-B$ negative examples to drive the network.
@@
$\mathcal{L}(y_{i}^{U},y_{j}^{L})=-\log{\frac{\exp{(\text{sim}_{tsp}(y_{i}^{U},y_{j}^{L})\times\tau)}}{\sum_{k=1}^{2B}{\mathbb{1}_{\left[k\neq i\right]}\exp{(\text{sim}_{tsp}(y_{i}^{U},y_{k}^{L})\times\tau)}}}}$ (3)
where $\mathbb{1}_{\left[k\neq j\right]}\in\{0,1\}$ is an indicator function evaluating to 1 if $k\neq i$. Finally, inspired by CLIP [4], we optimize a symmetric loss over these similarity scores within a mini-batch exhibited in Algorithm 1.
@@ 2.4 Simple Implementation
Algorithm 1 Pytorch-like pseudo-code for the core of an implement of OCL
The implementation of our OCL pre-training is efficient and involves minimal specialized operations. As pseudo code depicted in Algorithm 1, we make only minor modifications based on the MAE code, mainly involving the process subsequent to the acquisition of image feature embeddings from the encoder. First, we randomly mask a subset of embedded patch tokens with a low masking ratio. In succession, listed tokens are shuffled randomly and divided into two non-overlapping groups. Following MAE [2], within positional and [CLS] embeddings, lists of tokens are encoded by the ViT. It is noteworthy that we obtain the encoded [CLS] token from the ViT directly, without the incorporation of additional auxiliary modules, even a linear head or lightweight decoder.
@@ 3 Experiments
ViT-B/16 12 768 12 86M
ViT-L/16 24 1024 16 304M
ViT-h/16 32 1280 16 632M
Table 1: Configurations of Vision Transformer [10] models in our experiments. Block denotes the number of transformer blocks, with dim representing the input/output channel dimension of all blocks. Heads are the number of heads in multi-head attention modules. We also provide the parameter sizes of different models.
We perform self-supervised pre-training on the ImageNet-1K [17] dataset with the resolution of 224$\times$224. By default, ViT-B/16 and ViT-L/16 [10] are leveraged as the backbone architecture with 800 epochs for pre-training and 40 epochs for warm-up. ViT-B/16 applies the overall masked ratio of 0.3, while ViT-L/16 is set to 0.4. The initial base learning rate is $1.5\times 10^{-4}$. Like other contrastive learning [1, 4], our method relies on a large effective batch size: 9,600 for ViT-B/16 and 2,048 for ViT-L/16. Besides, $\kappa$ is set to 64 in the T-SP metric for computing similarity by default. Our implementation is based on MAE [2], with further details provided in the Supplementary Material.
In terms of supervised validation, OCL is evaluated through end-to-end fine-tuning and linear probing on the ImageNet-1k dataset for classification with 100 epochs for ViT-B/16 and 50 for ViT-L/16, following common practices [13, 2, 5]. Top-1 accuracy is utilized to verify the performance of different methods.